# Deterministic and Statistical Analysis of the DoF of Continuous Linear Arrays in the Near Field

### v1

@@ Abstract
This paper examines the number of communication modes, that is, the degrees of freedom (DoF), in a wireless setup comprising a small continuous linear intelligent antenna array in the near field of a large one. The framework allows for any orientations between the arrays and any positions in a two-dimensional space assuming that the transmitting array is placed at the origin. Therefore, apart from the length of the two continuous arrays, four key parameters determine the DoF and are hence considered in the analysis: the Cartesian coordinates of the center of the receiving array and two angles that model the rotation of each array around its center. The paper starts with the calculation of the deterministic DoF for a generic geometric setting, which extends beyond the widely studied paraxial case. Subsequently, a stochastic geometry framework is proposed to study the statistical DoF, as a first step towards the investigation of the system-level performance in near field networks. Numerical results applied to millimeter wave networks reveal the large number of DoF provided by near-field communications and unveiled key system-level insights.
@@ I Introduction
When two HMIMO devices communicate in far-field channels, the electromagnetic waves can be approximated as plane waves, and the system capacity depends on the richness of the scattering environment. In strong line-of-sight (LoS) channels, only one dominant path (i.e., with a strong coupling intensity) is available. In rich scattering environments, there are multiple strongly coupled communication modes (DoF), and appropriately designed orthogonal spatial beams form an optimal basis for the available spatial dimensions [14], [15], making beamspace MIMO communication possible [16, 17, 18, 19].
@@
The increasing interest in large intelligent surfaces (LIS) along with the relevance of higher frequency bands, such as the millimeter wave (mmWave) spectrum in current 5G systems and the sub-terahertz (THz) spectrum in future 6G systems, constitutes a significant change in communication paradigm, in which far-field communication models are inaccurate and need to be replaced by near-field channel models [20, 21]. Depending on the carrier frequency and the size of the LIS, the near-field region between two LISs may extend from tens to hundreds of meters. Consequently, unlike previous wireless generations, future LIS-aided 6G systems need to be designed based on near-field channel models where spherical waves replace the widely utilized plane waves [22]. This results in new opportunities for system design, such as the possibility of beam focusing in the near-field, in contrast to the conventional beam steering capability in far-field channels [23, 24, 25]. In [26], the authors derived the DoF when an LIS communicates with multiple single-antenna terminals in front of the LIS. The main result is that $2/\lambda$ terminals per meter can be spatially multiplexed for one-dimensional terminal deployments, while $\pi/\lambda^{2}$ terminals can be spatially multiplexed per square meter for two- and three-dimensional terminal deployments. The authors of [11] derived the DoF of a point-to-point HMIMO system with isotropic scattering. The main result is that the DoF is $2L_{x}/\lambda$ for one-dimensional linear arrays, $\pi L_{x}L_{y}/\lambda^{2}$ for two-dimensional planar arrays, and $2\pi L_{x}L_{y}/\lambda^{2}$ for three-dimensional volumetric arrays, where $L_{x},L_{y}$ are the lengths of the arrays. Recently, the authors of [27] applied the multidimensional sampling theorem and Fourier theory to study the Nyquist sampling and the DoF of an electromagnetic field under arbitrary scattering conditions. The main result is that the number of DoF per unit area is equal to the number of Nyquist samples per square meter to be able to reconstruct the field within a specified accuracy. The authors of [28] used the Fourier plane-wave series expansion for electromagnetic fields to evaluate the number of additional DoF in the reactive near-field, which is determined by the evanescent waves.
@@
The author of [29] introduced a general approach to evaluate the number of effective communication modes and their strength between two volumes. The solution is cast in terms of an eigenproblem obtained from the wave equation and Green’s function. An interesting result is that the number of DoF between two parallel communicating surfaces of area $A_{T}$ and $A_{R}$, located at a distance $r$, is equal to $A_{T}A_{R}/(\lambda^{2}r^{2})$, provided that the two surfaces are in the paraxial regime. Under the paraxial setting, to elaborate, $r$ is such that the Fresnel approximation can be applied, that is, only the first two terms of the Taylor series expansion of $r$ are retained, and the surface apertures are small compared to the propagation distance. Following a similar methodology, the authors of [30] derived a closed form expression for the DoF of LoS channels as a function of the orientations of the arrays, while still assuming the paraxial setting.
@@
Naturally, analysis of the DoF requires, especially in free-space channels, careful consideration of the transmission distance, the size of the surfaces, and the relative geometry of the surfaces. Thus, when the paraxial approximation, which is often assumed in the literature, is not fulfilled, that is, the transmission distance is comparable with the sizes of the surfaces, the calculation of the DoF is not straightforward, and alternative methods are required. More specifically, besides the numerical solution of the eigenproblem formulated in, e.g., [29], the DoF can be obtained by applying two general methods, which are known as (1) the spatial bandwidth [31]-[34] and (2) Landau’s methods [35]-[37], with the latter being highly coupled with the eigenproblem in [29]. Even though these methods are quite general and can be applied to general network scenarios, including non-paraxial settings, explicit analytical expressions for the DoF are difficult to obtain in general network deployments, e.g., in non-paraxial settings. Approximation approaches based on the spatial bandwidth and Landau’s eigenproblem, which can be applied to non-paraxial settings, can be found in [12] and [37], respectively. The impact of spatial blocking on the DoF has recently been analyzed in [38]. A short overview based on [4] and applied to continuous linear arrays can be found in [39]. To overcome these limitations, some ad hoc methods have recently been proposed.
For example, the authors of [40], have provided closed-form solutions for the DoF between a linear LIS and a linear small intelligent surface (SIS), which is of practical interest since the considered setting models the link between a base station and a mobile terminal. While the approach in [40] is a key step towards understanding non-paraxial settings, a limited geometric setup is considered, and the solutions are restricted to specific conditions.
@@
In the present work, a new analytical framework for calculating the DoF of a communication link between a large continuous linear array, and a small intelligent line, still modeled as a continuous linear array, is introduced. The proposed approach can be applied to paraxial and non-paraxial settings. To this end, the arrays are modeled as electromagnetic lines in free space, obeying the Huygens-Fresnel principle. The two arrays are assumed to radiate only in one of the two half-planes, in order to consider realistic deployments in which the transmitting array radiates only towards the intended receiving array. The transmitting array uses a steering function to concentrate the energy towards the direction of a focal point on the receiving array in the far-field. The calculation of the DoF is based on the calculation of the number of orthogonal steering functions that fit within the length of the receiving array. This approach allows us to account for near-field channels and non-paraxial settings. Therefore, the effective lengths of the two mutually visible arrays play an important role in the calculation of the DoF, since the mutual visibility of the arrays also determines the steering functionality. Another factor that determines the DoF is the projected lengths of the two linear arrays onto the line that is perpendicular to the line that connects the centers of the two arrays. These projections are highly dependent on the positions and orientations of the arrays.
@@
  • Using the methodology given in [40], we calculate the DoF for any geometric deployment of the arrays, considering the placement of the arrays anywhere in the horizontal plane and any orientation of the arrays. We also show that there exist optimal values for the rotating angles that maximize the DoF.
@@ II System Model
Since in the considered model the receiving array can be located anywhere in the plane and the two arrays may be rotated, there is the possibility that the receiving side of one array is not visible in its whole length from the transmitting side of the other array and vice versa. Two case studies with partial visibility are depicted in Fig. 2. Therefore, based on the values of $x_{0},y_{0},\theta_{T},\theta_{R}$, it is necessary to calculate the visible (effective) length of the two arrays $l_{T},l_{R}$, as well as the points of intersection of the arrays, denoted as $\eta_{i}$ and $\zeta_{i}$, respectively. Based on the location of these points, one may calculate the new centers $-L_{T}/2\leq\eta_{c}\leq L_{T}/2$, $-L_{R}/2\leq\zeta_{c}\leq L_{R}/2$, of the two effective array segments, respectively, and replace $\eta$ with $(\eta+\eta_{c})\in[-l_{T}/2,l_{T}/2]$ and $\zeta$ with $(\zeta+\zeta_{c})\in[-l_{R}/2,l_{R}/2]$ in (1), to account for the effective lengths.
@@
where $r(\eta)$ is defined in (1). If the Taylor series expansion is applied to $r(\eta)$ at $\eta=0$ and if one keeps only two terms of the series, then
$r(\eta)\approx r(0)+\eta\frac{\partial r(\eta)}{\partial\eta}\bigg{|}_{\eta=0}.$ (3)
The number of terms in the series to be kept is defined by the length of the transmitting array, when the variables $l_{R}$, $\theta_{T},\theta_{R}$, and $x_{0},y_{0}$ are fixed. When two terms are kept, the distance is approximated as a linear function of $\eta$. In detail, the first-order derivative in (3) is given by (4), shown at the bottom of this page.
@@
$r(\eta)\approx r(0)+\eta\,\rho=r(0)+\eta\,\frac{\sin\theta_{T}-\gamma\cos\theta_{T}}{\sqrt{1+\gamma^{2}}},$ (5)
@@
and $a$ is the angle formed by the segment from the center of the Tx, $\eta_{c}$, to the point $\zeta$ and the $x$ axis (see Fig. 1). According to the mean value theorem [45], the Lagrange remainder of Taylor’s series approximation of the function $r(\eta)$ at 0, is given by (7), shown at the bottom of the page,
$R_{1}(x)=\frac{r^{(2)}(c)}{2!}x^{2}=\frac{(x_{0}\cos\theta_{T}-\zeta\sin(\theta_{R}-\theta_{T})+y_{0}\sin\theta_{T})^{2}\,x^{2}}{2\,[(y_{0}+(\zeta+\zeta_{c})\cos\theta_{R}-(c+\eta_{c})\cos\theta_{T})^{2}+(x_{0}-(\zeta+\zeta_{c})\sin\theta_{R}+(c+\eta_{c})\sin\theta_{T})^{2}]^{3/2}}$ (7)
for $c\in(0,x)$. The full visibility condition corresponds to the case $\zeta_{c}=\eta_{c}=0$. Since (7) is increasing with $c$, and the maximum value for $x$ is $L_{T}/2$, the maximum error is obtained for $c=x=L_{T}/2$. By setting an upper bound on the error, one may compute the maximum transmitting array length $L_{T}$ for which the approximation is sufficiently accurate. If one drops the terms that are independent of $\eta$, which contribute with a constant phase shift to the focusing function, then (1) simplifies, using (5), to
$F_{T}(\eta)|_{\zeta}={\rm{rect}}\left(\frac{\eta}{l_{T}}\right)e^{j\frac{2\pi}{\lambda}\rho\eta}.$ (8)
In this special case, the approximation in (5) results in a linear phase as a function of $\eta$, resulting consequently in a beam steering phase profile. This means that the Tx concentrates energy towards the direction of the focal point in the far field and is not capable of focusing. However, since the Rx is a large array, the Tx is usually located in the near-field of the Rx, and a large number of modes are available. As detailed next, the proposed approach is also applicable to the near field, even though the approximation in (8) is based on a far-field model. Moreover, because of the reciprocity of the radio link, the DoF is the same if one switches the role of the two arrays from Tx/Rx to Rx/Tx. Therefore, there is no need to examine separately uplink and downlink transmissions, as far as the DoF is considered, as long as the definition of the orthonormal basis functions remains consistent.
@@ III Geometric Conditions For Visibility
III-A Basic Calculations
Based on Fig. 1, the endpoints of the two arrays are given by
@@ III-A Basic Calculations
The image of the functions in (11) are two lines between $T^{+}$ and $T^{-}$, and $R^{+}$ and $R^{-}$ respectively. If the domain, i.e., the parameters $\beta$ and $\delta$ are restricted to the interval $[0,1]$, the image is the segment between the endpoints. More precisely, if $\beta=0$, we obtain the point $T^{-}$, if $\beta=1$, we obtain the point $T^{+}$, and if $\beta=\frac{1}{2}$, we obtain the midpoint of the segment. The values $\beta<0$ correspond to the points of the line that extend beyond $T^{-}$, and the values $\beta>1$ correspond to the points that extend beyond $T^{+}$. The same applies to the parameter $\delta$.
@@
which will be used for a line-side check, especially when the Tx lies in the receiving side of the Rx and/or the Rx lies in the transmitting side of the Tx.
To this end, we introduce the magnitude of the cross product between two dimensional vectors as the determinant of the following $2\times 2$ matrix:
@@
Specifically, if $\mathbf{r}\times(-\mathbf{c})>0$, $T^{0}$ is located to the left of Rx and within its receiving half plane; if $\mathbf{r}\times(-\mathbf{c})<0$ then $T^{0}$ is located to the right of the Rx and on the other half plane; and if $\mathbf{r}\times(-\mathbf{c})=0$, $T^{0}$ is located on the Rx array.
The second line-side check is performed for the $R^{0}$ point with respect to the Tx, by using the cross-product
@@
and similar conclusions are for $T^{0}$ can be drawn.
The vectors normal to $\mathbf{t}$ and $\mathbf{r}$ that are directed toward the transmitting and receiving side of the two arrays, respectively, are called inward edge normal vectors and are calculated using the cross product of two dimensional vectors
@@
The coordinates $(P_{x},P_{y})$ of the point $P$, given by (20), shown at the bottom of the next page, are calculated by substituting the solution for the parameter $\delta_{P}$ for the line $r$ or the parameter $\beta_{P}$ for the line $t$. The parameter $\zeta_{i}$ (the distance between $P$ and $R^{0}$) on the line $r$ of the Rx array is given by
@@
Therefore, the new centers of the Tx and Rx arrays can be calculated when the point of intersection lies on their line segments, i.e., when $0\leq\beta_{P}\leq 1$ or $0\leq\delta_{P}\leq 1$, and the parameters $\zeta_{i}$ and $\eta_{i}$ take values in the range $[-L_{R}/2,L_{R}/2]$, $[-L_{T}/2,L_{T}/2]$, respectively. The coordinates of the new center on the Rx array when the endpoint $R^{-}$ is visible from the Tx array are given by
Following the same procedure for the Tx array, the distance of the new center from the origin is given by replacing $L_{R}$ with $L_{T}$, $\zeta_{i}$ with $\eta_{i}$, and $R^{-},R^{+}$ with $T^{-},T^{+}$ in (23).
Finally, the new lengths of the arrays are given by
and
III-B Proposed Visibility Algorithm
Based on these preliminary calculations, the algorithmic steps to evaluate the visibility state between the two arrays are given in Algorithm 1. In the following, four auxiliary vectors are defined: $\mathbf{t_{-}}=(T^{-}-R^{0})$, $\mathbf{t_{+}}=(T^{+}-R^{0})$, $\mathbf{r_{-}}=(R^{-}-T^{0})$, and $\mathbf{r_{-}}=R^{-}-T^{0}$. The algorithm utilizes the cross products $\mathbf{r}\times(-\mathbf{c})$ and $\mathbf{t}\times\mathbf{c}$, given by (14) and (15), respectively, as well as the following magnitudes of cross products:

### v2

@@ Abstract
This paper examines the number of communication modes, that is, the degrees of freedom (DoF) in a wireless line-of-sight channel comprising a small continuous linear intelligent antenna array in the near field of a large one. The framework allows for any orientations between the arrays and any positions in a two-dimensional space assuming that the transmitting array is placed at the origin. Therefore, apart from the length of the two continuous arrays, four key parameters determine the DoF and are hence considered in the analysis: the Cartesian coordinates of the center of the receiving array and two angles that model the rotation of each array around its center. The paper starts with the calculation of the deterministic DoF for a generic geometric setting, which extends beyond the widely studied paraxial case. Subsequently, a stochastic geometry framework is proposed to study the statistical DoF, as a first step towards the investigation of the system-level performance in near field networks. Numerical results applied to millimeter wave networks reveal the large number of DoF provided by near-field communications and unveil key system-level insights. A comparison of the proposed method with the singular value decomposition-based method is illustrated to validate the model.
@@ I Introduction
When two HMIMO devices communicate in far-field channels, the electromagnetic waves can be approximated as plane waves, and the system capacity depends on the richness of the scattering environment. In strong line-of-sight (LoS) channels, only one dominant path (i.e., with a strong coupling intensity) is available. In rich scattering environments, there are multiple strongly coupled communication modes (DoF), and appropriately designed orthogonal spatial beams form an optimal basis for the available spatial dimensions [14], [15], making beamspace MIMO communication possible [16, 17, 18, 19]. Nevertheless, when considering wave propagation in intricate scattering environments, such as urban settings, where the wavelength is significantly shorter than the typical size of the scatterers, the wave scattering process may exhibit chaotic ray dynamics. This phenomenon necessitates the use of a stochastic Green’s function [20], which serves as the probabilistic solution to the wave equation in chaotic media. This approach is grounded in random matrix theory. Another approach is presented in [21], where the authors developed a Rayleigh fading model. In this model, the channel impulse response is a spatially stationary circularly symmetric complex Gaussian electromagnetic random field. The authors obtained a Fourier spectral representation of the random channel that accurately describes the impulse response only asymptotically, i.e., as the normalized array size approaches infinity.
@@
The increasing interest in large intelligent surfaces (LIS) along with the relevance of higher frequency bands, such as the millimeter wave (mmWave) spectrum in current 5G systems and the sub-terahertz (THz) spectrum in future 6G systems, constitutes a significant change in communication paradigm, in which far-field communication models are inaccurate and need to be replaced by near-field channel models [22, 23]. Depending on the carrier frequency and the size of the LIS, the near-field region between two LISs may extend from tens to several hundreds of meters. Consequently, unlike previous wireless generations, future LIS-aided 6G systems need to be designed based on near-field channel models where spherical waves replace the widely utilized plane waves [24]. This results in new opportunities for system design, such as the possibility of beam focusing in the near-field, in contrast to the conventional beam steering capability in far-field channels [25, 26, 27]. In [28], the authors derived the DoF when an LIS communicates with multiple single-antenna terminals in front of the LIS. The main result is that $2/\lambda$ terminals per meter can be spatially multiplexed for 1-D terminal deployments, while $\pi/\lambda^{2}$ terminals can be spatially multiplexed per square meter for 2-D and 3-D terminal deployments. The authors of [11] derived the DoF of a point-to-point HMIMO system with isotropic scattering. The main result is that the DoF is $2L_{x}/\lambda$ for 1-D linear arrays, $\pi L_{x}L_{y}/\lambda^{2}$ for 2-D planar arrays, and $2\pi L_{x}L_{y}/\lambda^{2}$ for 3-D volumetric arrays, where $L_{x},L_{y}$ are the lengths of the arrays. Recently, the authors of [29] applied the multidimensional sampling theorem and Fourier theory to study the Nyquist sampling and the DoF of an electromagnetic field under arbitrary scattering conditions. The main result is that the number of DoF per unit area is equal to the number of Nyquist samples per square meter to be able to reconstruct the field within a specified accuracy. The authors of [30] used the Fourier plane-wave series expansion for electromagnetic fields to evaluate the number of additional DoF in the reactive near-field, which is determined by the evanescent waves.
@@
The author of [31] introduced a general approach to evaluate the number of effective communication modes and their strength between two volumes. The solution is cast in terms of an eigenproblem obtained from the wave equation and Green’s function. An interesting result is that a heuristic number of well-coupled modes (DoF) between two parallel communicating surfaces of area $A_{T}$ and $A_{R}$, located at a distance $r$, is equal to $A_{T}A_{R}/(\lambda^{2}r^{2})$, provided that the two surfaces are in the paraxial regime. Under the paraxial setting, to elaborate, $r$ is such that the Fresnel approximation can be applied, that is, only the first two terms of the Taylor series expansion of $r$ are retained, and the surface apertures are small compared to the propagation distance. This implies that the spherical wavefront is approximated by a parabola, leading to the term parabolic wavefront model [32]. Following a similar methodology, the authors of [33] derived a closed form expression for the DoF of LoS channels as a function of the orientations of the arrays, while still assuming the paraxial setting.
@@
Naturally, analysis of the DoF requires, especially in free-space channels, careful consideration of the transmission distance, the size of the surfaces, and the relative geometry of the surfaces. Thus, when the paraxial approximation, which is often assumed in the literature, is not fulfilled, that is, the transmission distance is comparable with the sizes of the surfaces, the calculation of the DoF is not straightforward, and alternative methods are required. More specifically, besides the numerical solution of the eigenproblem formulated in, e.g., [31], the DoF can be obtained by applying two general methods, which are known as (1) the spatial bandwidth [34]-[37] and (2) Landau’s methods [38]-[40], with the latter being highly coupled with the eigenproblem in [31]. Even though these methods are quite general and can be applied to general network scenarios, including non-paraxial settings, explicit analytical expressions for the DoF are difficult to obtain in general network deployments, e.g., in non-paraxial settings. Approximation approaches based on the spatial bandwidth and Landau’s eigenproblem, which can be applied to non-paraxial settings, can be found in [12] and [40], respectively. The impact of spatial blocking on the DoF has recently been analyzed in [41]. A short overview based on [4] and applied to continuous linear arrays can be found in [42]. To overcome these limitations, some ad hoc methods have recently been proposed. For example, the authors of [43], have provided closed-form solutions for the DoF between a linear LIS and a linear small intelligent surface (SIS), which is of practical interest since the considered setting models the link between a base station and a mobile terminal. While the approach in [43] is a key step towards understanding non-paraxial settings, a limited geometric setup is considered, and the solutions are restricted to specific conditions.
@@
In this present work, a new analytical framework is introduced to calculate the number of DoFs of a communication link between two coplanar continuous linear arrays, a large and a small one, under LoS channel conditions. The proposed approach can be applied to paraxial and non-paraxial settings. To this end, the arrays are modeled as electromagnetic lines in free space, obeying the Huygens-Fresnel principle. The two arrays are assumed to radiate only in one of the two half-planes, in order to consider realistic deployments in which the transmitting array radiates only towards the intended receiving array. The transmitting array uses a focusing function to concentrate the energy in the direction of a focal point on the receiving array. The calculation of the DoF is based on the calculation of the number of orthogonal focusing functions that fit within the length of the receiving array. This approach allows us to account for near-field channels and non-paraxial settings. Therefore, the effective lengths of the two mutually visible arrays play an important role in the calculation of the DoF, since the mutual visibility of the arrays also determines the focusing functionality. Another factor that determines the DoF is the projected lengths of the two linear arrays onto the line that is perpendicular to the line that connects the centers of the two arrays. These projections are highly dependent on the positions and orientations of the arrays.
@@
  • We extend the methodology given in [43] by introducing a Taylor series expansion of the distance $r$ with one more term and derive the corresponding focusing function. Then the DoF are calculated for almost any geometric deployment of the arrays, considering the placement of the arrays almost anywhere on the horizontal plane and any orientation of the arrays.
  • We provide insights into the dependence of the number of DoF as a function of the geometrical parameters, that is, on the coordinates of the centers of the two arrays, their effective lengths, and the rotation angles. Comprehensive numerical results for a wide range of paraxial and non-paraxial settings are presented.
  • We validate the proposed method to compute the number of DoF against the SVD numerical method, by considering free space channels.
@@ II System Model
Since in the considered model the receiving array can be located anywhere on the plane and the two arrays may be rotated, there is the possibility that the receiving side of one array is not visible in its whole length from the transmitting side of the other array and vice versa. Two case studies with partial visibility are depicted in Fig. 2. Therefore, based on the values of $x_{0},y_{0},\theta_{T},\theta_{R}$, it is necessary to calculate the visible (effective) length of the two arrays $l_{T},l_{R}$, as well as the points of intersection of the arrays, denoted as $\eta_{i}$ and $\zeta_{i}$, respectively. Based on the location of these points, one may calculate the new centers $-L_{T}/2\leq\eta_{c}\leq L_{T}/2$, $-L_{R}/2\leq\zeta_{c}\leq L_{R}/2$, of the two effective array segments, respectively, and replace $\eta$ with $(\eta+\eta_{c})$ and $\zeta$ with $(\zeta+\zeta_{c})$ in (1), to account for the effective lengths. Moreover, $(\eta-\eta_{c})\in[-l_{T}/2,l_{T}/2]$ and $(\zeta-\zeta_{c})\in[-l_{R}/2,l_{R}/2]$.
@@
where $r(\eta)$ is defined in (1). If the Taylor series expansion is applied to $r(\eta)$ at $\eta=0$ and if one keeps the first three terms of the series, then
$r(\eta)\approx r(0)+\eta\frac{\partial r(\eta)}{\partial\eta}\bigg{|}_{\eta=0}+\frac{\eta^{2}}{2}\frac{\partial^{2}r(\eta)}{\partial\eta^{2}}\bigg{|}_{\eta=0}.$ (3)
In detail, the first- and second-order derivatives in (3) are given by (4) and (5), shown at the bottom of this page.
@@
$\frac{\partial^{2}r(\eta)}{\partial\eta^{2}}=\frac{(x_{0}\cos\theta_{T}-(\zeta+\zeta_{c})\sin(\theta_{R}-\theta_{T})+y_{0}\sin\theta_{T})^{2}}{[(y_{0}+(\zeta+\zeta_{c})\cos\theta_{R}-(\eta+\eta_{c})\cos\theta_{T})^{2}+(x_{0}-(\zeta+\zeta_{c})\sin\theta_{R}+(\eta+\eta_{c})\sin\theta_{T})^{2}]^{3/2}}$ (5)
@@
$r(\eta)\approx r(0)+\eta\,\rho+\eta^{2}\,\tilde{\rho},$ (6)
@@
$\rho=\frac{\sin\theta_{T}-\gamma\cos\theta_{T}}{\sqrt{1+\gamma^{2}}},$ (7)
$\tilde{\rho}$ $=(x_{0}-(\zeta+\zeta_{c})\sin\theta_{R}+\eta_{c}\sin\theta_{T})^{-\frac{5}{2}}\frac{(\cos\theta_{T}+\gamma\sin\theta_{T})^{2}}{2(1+\gamma^{2})^{3/2}}$ (8)
$=(x_{0}-(\zeta+\zeta_{c})\sin\theta_{R}+\eta_{c}\sin\theta_{T})^{-\frac{5}{2}}\frac{(1+\gamma^{2})^{\frac{3}{2}}}{2}\left(\frac{\partial\rho}{\partial\gamma}\right)^{2},$
@@
and $a$ is the angle formed by the segment from the center of the Tx, $\eta_{c}$, to the point $\zeta$ and the $x$ axis (see Fig. 1). It is noted that the expansion in (3) is different from the expansion given in [32] for the parabolic wavefront approximation, $r\approx r_{x}+\frac{r_{y}^{2}}{2d_{0}}$, where $d_{0}=\sqrt{x_{0}^{2}+y_{0}^{2}}$, providing a considerably reduced approximation error in non-paraxial settings.
If one drops the terms that are independent of $\eta$, which contribute with a constant phase shift to the focusing function, then (2) simplifies, using (6), to
$F_{T}(\eta)|_{\zeta}={\rm{rect}}\left(\frac{\eta}{l_{T}}\right)e^{j\frac{2\pi}{\lambda}(\rho\,\eta+\eta^{2}\tilde{\rho})}.$ (10)
Therefore, the approximation in (6) results in a quadratic phase as a function of $\eta$, which results in a focusing function in the near field. The requirement for a quadratic polynomial originates from the fact that the boresight Fraunhofer distance for a Tx linear array of length $L_{T}$ and an Rx linear array of length $L_{R}$ is $d^{ff}=\frac{2(L_{T}+L_{R})^{2}}{\lambda}$, [23], [47]. Therefore, the Tx concentrates energy in the direction of the focal point on the Rx array. The distance of the two arrays also determines the depth of the focusing beam. If the Rx array is placed in the far field, then the focusing function degenerates into a beam steering phase profile, resulting in an infinite depth beamforming. Because of the reciprocity of the radio link, the DoF is the same if one switches the role of the two arrays from Tx/Rx to Rx/Tx. Therefore, there is no need to examine separately uplink and downlink transmissions, as far as the DoF is considered, as long as the definition of the orthonormal basis functions remains consistent.
@@ III Geometric Conditions For Visibility
III-A Description of the Visibility Algorithm
This subsection presents the main methodological steps to identify the visibility condition between the two arrays, which corresponds to identifying the case, among the four discussed, to be considered, as well as to compute the parameters required for the calculation of the number of DoF. The visibility condition is governed by the length and orientation of the two arrays. For this reason, two segments are defined using the endpoints of the arrays and two lines in parametric form. Based on Fig. 1, the endpoints of the two arrays are given by
@@ III-A Description of the Visibility Algorithm
The image of the functions in (13) are two lines between $T^{+}$ and $T^{-}$, and $R^{+}$ and $R^{-}$ respectively. If the domain, i.e., the parameters $\beta$ and $\delta$ are restricted to the interval $[0,1]$, the image is the segment between the endpoints. More precisely, if $\beta=0$, we obtain the point $T^{-}$, if $\beta=1$, we obtain the point $T^{+}$, and if $\beta=\frac{1}{2}$, we obtain the midpoint of the segment. The values $\beta<0$ correspond to the points of the line that extend beyond $T^{-}$, and the values $\beta>1$ correspond to the points that extend beyond $T^{+}$. The same applies to the parameter $\delta$. The two lines may intersect, and in such a case, the point of intersection needs to be calculated and subsequently verified to determine whether it falls within the Tx or Rx array. Additionally, the algorithm applies a line-side check, which entails determining whether the Rx array is positioned to the left or right of the Tx and, further, whether the receiving side of the Rx array faces the transmitting side of the Tx array. For these computations, a computer graphics approach is employed and additional vector definitions are introduced next, including the outer product for two-dimensional vectors.
@@
which will be used for a line-side check, especially when the Tx lies in the receiving side of the Rx and/or the Rx lies in the transmitting side of the Tx. To this end, we introduce the magnitude of the cross product between two dimensional vectors as the determinant of the following $2\times 2$ matrix:
@@
Specifically, if $\mathbf{r}\times(-\mathbf{c})>0$, $T^{0}$ is located to the left of Rx and within its receiving half plane; if $\mathbf{r}\times(-\mathbf{c})<0$ then $T^{0}$ is located to the right of the Rx and on the other half plane; and if $\mathbf{r}\times(-\mathbf{c})=0$, $T^{0}$ is located on the Rx array. The second line-side check is performed for the $R^{0}$ point with respect to the Tx, by using the cross-product
@@
and similar conclusions for $T^{0}$ can be drawn. The vectors normal to $\mathbf{t}$ and $\mathbf{r}$ that are directed toward the transmitting and receiving side of the two arrays, respectively, are called inward edge normal vectors and are calculated using the cross product of two dimensional vectors
@@
The coordinates $(P_{x},P_{y})$ of the point $P$, given by (23), shown at the bottom of the page, are calculated by substituting the solution for the parameter $\delta_{P}$ for the line $r$ or the parameter $\beta_{P}$ for the line $t$. The parameter $\zeta_{i}$ (the distance between $P$ and $R^{0}$) on the line $r$ of the Rx array is given by
@@
Therefore, the new centers of the Tx and Rx arrays can be calculated when the point of intersection lies on their line segments, i.e., when $0\leq\beta_{P}\leq 1$ or $0\leq\delta_{P}\leq 1$, and the parameters $\zeta_{i}$ and $\eta_{i}$ take values in the range $[-L_{R}/2,L_{R}/2]$, $[-L_{T}/2,L_{T}/2]$, respectively. Therefore, in partial visibility conditions, to calculate the effective lengths of the arrays, one has to differentiate between four cases: i) partial Tx visibility and the point $T^{-}$ is visible from the Rx array, ii) partial Tx visibility and the point $T^{+}$ is visible, iii) partial Rx visibility and the point $R^{-}$ is visible from the Tx array, and iv) partial Rx visibility and the point $R^{+}$ is visible. This is accomplished by defining four auxiliary vectors: $\mathbf{t_{-}}=(T^{-}-R^{0})$, $\mathbf{t_{+}}=(T^{+}-R^{0})$, $\mathbf{r_{-}}=(R^{-}-T^{0})$, and $\mathbf{r_{-}}=R^{-}-T^{0}$, and the following magnitudes of cross products:
@@
. The coordinates of the new center on the Rx array when the endpoint $R^{-}$ is visible from the Tx array are given by
Following the same procedure for the Tx array, the distance of the new center from the origin is given by replacing $L_{R}$ with $L_{T}$, $\zeta_{i}$ with $\eta_{i}$, and $R^{-},R^{+}$ with $T^{-},T^{+}$ in (29). Finally, the new lengths of the arrays are given by
and
III-B Proposed Visibility Algorithm
Based on the structure and definitions presented in the preceding subsection, the algorithmic steps to assess the visibility condition between the two arrays and to compute the new centers and effective lengths are provided in Algorithm 1.